# A new expansion of the Riemann zeta function

### v1

@@ Abstract
After a brief introduction to Ramanujan’s method of summation, we give an expansion of the Riemann Zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$ where the functions $P_{m}$ are polynomials with their roots on the line $\{\Re(s)=1/2\}$, the coefficients $x_{m}$ being finite linear combinations of the Euler constant $\gamma$ and the values $\zeta(2),\zeta(3),\dots,\zeta(m+1).$
@@ 1 Introduction
There are many polynomial expansions of the Riemann zeta function. Some of them (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta. For example
@@
Others are linked to the well known expansion
@@
where $\gamma_{m}$ are the Stieltjes constants. This gives (cf.[E-G])
$\zeta(s)-\frac{s}{s-1}=\sum_{m=0}^{+\infty}(-1)^{m}l_{m}(\frac{s-1}{s})^{m},$
with
$l_{m}=\sum_{k=1}^{m}\binom{m-1}{k-1}(-1)^{m-k}\frac{\gamma_{k}}{k!}.$
and for $s=\frac{1}{2}+it$ $(t\in\mathbb{R}$) one obtain an expansion of $\zeta(\frac{1}{2}+it)$ on the orthonormal basis $(e_{m})$ of $L^{2}(\mathbb{R},\frac{dt}{2\pi(t^{2}+1/4)})$ given by
$e_{m}(t)=\Big(\frac{1-2it}{1+2it}\Big)^{m}.$
There is also (cf. [K]) an expansion of the function
$\Xi(t)=\frac{s(s-1)}{2}\pi^{-s/2}\Gamma(s/2)\zeta(s),\ \ \ s=\frac{1}{2}+it,\ (t\in\mathbb{R}),$
which takes the form
$\Xi(t)=e^{\pi t/4}\sum_{m=0}^{+\infty}b_{m}p_{m}(t),$
where $p_{m}(t)$ are polynomials related to the Mellin transform of odd Hermite functions. These polynomials are orthogonal in $(L^{2}(\mathbb{R},|\Gamma(3/4+it/2)|^{2}dt)$, but it seems that the coefficients $b_{m}$ does not have a simple form.
The polynomial expansion we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ are the Laguerre polynomials defined by
@@
$e^{-xz}=\sum_{m=0}^{+\infty}2e^{-x}L_{m}(2x)\psi_{m}(z).$
where $(\psi_{m})_{m\geq 0}$ the sequence of the functions defined in the half-plane $\{Re(z)>0\}$ by
$\psi_{m}(z)=\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m}.$
@@
$\Gamma(s)z^{-s}=\sum_{m=0}^{+\infty}\Big(2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx\Big)\psi_{m}(z),$
Since we can write
$2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx={\Gamma(s)}P_{m}(s),$
where $P_{m}$ is the polynomial
$P_{m}(s)=\sum_{k=0}^{m}\binom{m}{k}(-1)^{k}2^{k+1}\ \frac{\Gamma(s+k)}{\Gamma(s)}\frac{1}{k!},$
then we have, for $\Re(s)>0$, the expansion
$z^{-s}=\sum_{m=0}^{+\infty}\psi_{m}(z)P_{m}(s).$
@@
$\sum_{n\geq 1}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}\Big(\sum_{n\geq 1}\psi_{m}(n)\Big)P_{m}(s).$
In section 3, we prove that, for $0<\Re(s)<1$, this last formula is correct if we replace the divergent series $\sum_{n\geq 1}$ by their Ramanujan’s summations $\sum_{n\geq 1}^{\mathcal{R}}$, and then we have
@@
In section 2 we give a short introduction to Ramanujan’s summation of divergent series. Since we will see that, for $s\neq 1$, we have
@@
and $r_{m}$ are rational numbers explicitely given. For $s=\frac{1}{2}+it$ with $t\in\mathbb{R}$, the preceding relation gives the expansion
@@
where the polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$
form complete orthonormal system in $\big(L^{2}(-\infty,+\infty),\frac{dt}{\cosh(\pi t)}\big)$. Thus, for every $m\geq 1$, the roots of $Q_{m}$ are real, and this implies that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$.
@@ 2.1 Ramanujan’s constant of a series
Let a function $f\in C^{1}(]0,+\infty[)$, the first order Euler-MacLaurin formula says that, for every integer $n\geq 1$
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+\frac{1}{2}\big(f(n)+f(1)\big)+\int_{1}^{n}\big(x-[x]-\frac{1}{2}\big)f^{\prime}(x)\ dx$}, (1)
where $[x]$ denote the integer part of $x$. For $x\geq 1$, let
$b(x)=x-[x]-\frac{1}{2}$
and let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
@@
We observe that $C(f)$ is also well defined for some divergent series. We have immediately $C(1)=1/2.$ And applying (2) to the function $f:x\mapsto 1/x$, we have for every integer $n\geq 1$
@@
which gives that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
$C(f)=\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
@@ 2.2 Plana’s formula
Let $f$ be an analytic function in the half-plane $\{\Re(z)>0\}$ such that there is a constant ${a}<2\pi$ and, for any interval $c,d]\subset]0,+\infty[$, a constant $C_{c,d}>0$ with
$\left|f(x+it)\right|\leq C_{c,d}\ e^{a|t|},\ \text{ for }x\in[c,d]\ \text{ and all }t\in\mathbb{R}.$
@@
If we now compare Euler-MacLaurin formula (1) and Plana’s (3), then we get, for $n\geq 1$,
@@
If the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\,dx$ is convergent and if
@@
Thus, in this case the Ramanujan’s constant $C(f)$, previously defined by (2), is also
@@
Take now this last formula as the definition of Ramanujan’s constant $C(f)$ for a function $f$ verifying the hypotheses of theorem 1.
@@ 2.3 The space $\mathcal{E}^{\pi}$
Let ${\alpha}>0$ and ${\mathcal{E}}^{\alpha}$ the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\alpha$ and $\emph{b}>0$ with :
@@
Thus $f_{s}\in\mathcal{E}^{\alpha}$ if and only if $\alpha<\Im(s)<\alpha$.
Properties
a) If $f\in\mathcal{E}^{\alpha}$ and $g\in\mathcal{E}^{\beta}$, alors $fg\in E^{\alpha+\beta}$.
b) If $f\in\mathcal{E}^{\alpha}$ and if $F(z)=\int_{1}^{z}f(u)du$ then $F\in\mathcal{E}^{\alpha}$.
@@
If $f\in\mathcal{E}^{\pi}$ is such that $f(n)=0$ for all $n\;=\;1,\,2,\,\ldots$, then $f\equiv 0$.
Thus a function $f\in\mathcal{E}^{\pi}$ is only determined by the values $f(1),f(2),f(3),\dots$.
@@
The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}^{\pi}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
@@ 2.4 Ramanujan’s method of summation
Let $f$ and $g$ two functions verifying the hypothesis of theorem 1. By Carlson’s theorem we know that if $f\in\mathcal{E}^{\pi}$ and $g\in\mathcal{E}^{\pi}$ verify f(n)=g(n){f(n)=g(n)} for $n=1,2,3,\dots$, then for x>0{x>0} and $t\in\mathbb{R}$, we have
@@
For $f\in\mathcal{E}^{\pi}$, we define the Ramanujan’s summation of the series $\sum_{n\geq 1}f(n)$ by
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=\frac{f(1)}{2}+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt\,.$ (6)
@@
Let $f\in\mathcal{E}^{\pi}$ then by (3), we have for every integer $n\geq 1$
@@
and if in addition, the integral $\int_{1}^{+\infty}f(u)\ du$ is convergent then the series $\sum_{n\geq 1}f(n)$ is convergent. In this case we have by (7)
@@
For $f:x\mapsto\frac{1}{x^{s}}$, with $\Re(s)>1$, we get
@@ 2.5 Interpolation of partial sums
To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of partial sums of this series . In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
@@
a) Existence. Since $f\in\mathcal{E}^{\pi}$ then the function $\varphi_{f}$, defined for $\Re(z)>0$, by
@@
is also in $\mathcal{E}^{\pi}$. By (3) we have $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)$ for every integer $n\geq 1$, thus the function $\varphi_{f}$ verifies
@@
Since the functions $z\mapsto\varphi_{f}(z+1)-\varphi_{f}(z)$ and $z\mapsto f(z+1)$ are in $\mathcal{E}^{\pi}$, then by Carlson’s theorem, we get
@@
The condition $\varphi_{f}(1)=f(1)$ is simply a consequence of (6) and (10).
b) Uniqueness. If $\varphi\in\mathcal{E}^{\pi}$ is such that
@@
thus $\varphi(n)=\sum_{k=1}^{n}f(k)=\varphi_{f}(n)$ for every integer $n\geq 1$. Since $\varphi$ and $\varphi_{f}$ are in the space $\mathcal{E}^{\pi}$, then by Carlson’s theorem we get $\varphi=\varphi_{f}.$
@@
The preceding proof asserts that if $f\in\mathcal{E}^{\pi}$ then
@@
The function $\varphi_{f}$ given by (10) is the unique function in $\mathcal{E}^{\pi}$ wich interpolates the partial sums of the series $\sum_{n\geq 1}f(n)$. Note that, by (11), we have for $\Re(z)>0$
@@
Let $f:z\mapsto 1/z$, then $\varphi_{f}$ is related to digamma function which is defined, for $\Re(z)>0$, by
@@
This function is analytic in $\{\Re(z)>0\}$, and by logarithmic differentiation of the classical product formula for $\Gamma$ (cf. [G-R]) we get for $\Re(z)>0$
$\Psi(z)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\big(\frac{1}{n}-\frac{1}{n+z}\big)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\frac{z}{n(n+z)}$ (12)
Then it is easily verified that $\Psi\in\mathcal{E}^{\pi}$ and that
$\Psi(z+2)\;-\;\Psi(z+1)=\frac{1}{z+1}$
@@
A very useful formula for Ramanujan’s summation is provided by the following theorem.
@@
Let $f\in\mathcal{E}^{\pi}$, and let $\varphi\in\mathcal{E}^{\pi}$ be a solution of
$\varphi(x+1)-\varphi(x)$ $=$ $f(x+1),\ \text{ for all }\ x>0,$
$\varphi(1)$ $=$ $f(1).$
@@
$\int_{1}^{2}\varphi_{f}(x)\ dx\;\;=\int_{1}^{2}F(u)\ du\,+\,\frac{1}{2}\int_{1}^{2}f(x)\ dx\,+\sum_{n\geq 1}^{\mathcal{R}}f(n)-iD(f),$ (13)
@@
If we apply Plana’s formula to the function $F$, with $n=2$, then we get
$-\big(\frac{1}{2}F(1)+\frac{1}{2}F(2)\big)+\!\int_{1}^{2}F(t)dt$ $=i\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}dt$
$-i\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}dt,$
which is
@@
Let $f\in\mathcal{E}^{\pi}$, then the function $x\mapsto f(x+1)$ belongs to $\mathcal{E}^{\pi}$ and we have the shift formula
$\sum_{n\geq 1}^{\mathcal{R}}f(n+1)\;\;=\sum_{n\geq 1}^{\mathcal{R}}f(n)-f(1)+\int_{0}^{1}f(x+1)\ dx.$
@@
Let $f:x\mapsto 1/x$. By example 3, we have $\varphi_{f}(x)=\Psi(x+1)+\gamma,$ thus we get
@@
Observe that, by corollary 1, we have
@@
Since, fro any $x>0$, we have
$\log(\Gamma(x+2))-\log(\Gamma(x+1))=\log(x+1)$
@@
Let $f(x)=\Psi(x)$ and $\varphi(x)=x\Psi(x)-x+1$. We have $f(1)=\Psi(1)=\varphi(1)$ and
@@
Thus, by the previous theorem, we get
@@
then by Raabe’s formula
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=-\log(\sqrt{2\pi})+\frac{1}{2}.$
@@
We can evaluate $\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ for $k=1,2,3,\dots$ by means of Bernoulli numbers. Let the Bernoulli polynomials $B_{n}(x)$ be defined by $B_{0}(x)=1$ and, for $k\geq 1,$
$B^{\prime}_{k}(x)=nB_{k-1}(x)\ \text{ with }\ \int_{0}^{1}B_{k}(x)dx=0.$
The Bernoulli numbers are the numbers $B_{k}=B_{k}(0)$. These polynomials verify (cf. [G-R])
@@
For $k\geq 1$, we have $B_{k+1}(1)=B_{k+1}$. Let the function $f:x\mapsto x^{k}$ with an integer $k\geq 1$, then by theorem 3, we obtain
@@
$\sum_{n\geq 1}^{\mathcal{R}}n^{0}=\frac{1}{2}=\zeta(0)+1,$
and by (9) we have in the half-plane $\{\Re(s)>1\}$
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\zeta(s)-\frac{1}{s-1}.$ (16)
All these formulae are special cases the following general result.
@@
$\zeta(s)=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}+\frac{1}{s-1}$ (17)
@@
First, it is easily seen that, for every $s\in\mathbb{C}$, the function $z\mapsto 1/{z^{s}}$ is in the space $\mathcal{E}^{\alpha}$, for all $\alpha>0.$ Since, if $s=\sigma+i\tau$ ($\sigma\in\mathbb{R},\ \tau\in\mathbb{R}$), we can write
@@
where $z=x+it$, with $x>\varepsilon>0$ and $t\in\mathbb{R}$. Thus the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is well defined and, by (6), we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\;\;\frac{1}{2}+i\int_{0}^{+\infty}\Big({\frac{1}{(1+it)^{s}}-\frac{1}{(1-it)^{s}}}\big)\frac{1}{e^{2\pi t}-1}\,\ dt\,$

### v2

@@ Abstract
We introduce a family of orthogonal polynomials $P_{m}(s)$, whose roots lie on the line $\{\Re(s)=1/2\}$, and we give an expansion of the Riemann zeta function in the critical strip as a convergent series $\sum_{m\geq 0}x_{m}P_{m}(s)$. The coefficients $x_{m}$of this expansion are expressed as linear combinations with rational coefficients of the Euler constant $\gamma$ and the values $\zeta(2),$ $\zeta(3),$ $\dots,$ $\zeta(m+1).$
@@ 1 Introduction
Some polynomial expansions of the Riemann zeta function (cf.[E-K-K], [F-V],[B],[M], [V]) are based on the Newton interpolation formula for functions related to Zeta. For example
@@
Others (cf.[E-G]) are linked to the well known expansion
@@
where $\gamma_{m}$ are the Stieltjes constants.
The polynomial expansion of the Riemann zeta function that we give in this paper is related to the Mellin transform of the functions $x\mapsto e^{-x}L_{m}(2x)$ where $(L_{m})_{m\geq 0}$ is the sequence of Laguerre polynomials defined by
@@
$e^{-xz}=\sum_{m=0}^{+\infty}2e^{-x}L_{m}(2x)\big(\frac{z-1}{z+1}\big)^{m}\frac{1}{1+z}.$
@@
$z^{-s}=\sum_{m=0}^{+\infty}\frac{1}{1+z}\big(\frac{z-1}{z+1}\big)^{m}P_{m}(s),$
where $P_{m}$ is a polynomial defined by
${\Gamma(s)}P_{m}(s)=2\int_{0}^{+\infty}e^{-x}L_{m}(2x)x^{s-1}\ dx.$
@@
$\sum_{n\geq 1}\frac{1}{n^{s}}=\sum_{m=0}^{+\infty}\Big(\sum_{n\geq 1}\frac{1}{1+n}\big(\frac{n-1}{n+1}\big)^{m}\Big)P_{m}(s).$
We prove that, for $0<\Re(s)<1$, this last formula is correct if we replace the divergent series $\sum_{n\geq 1}$ by their Ramanujan’s summations $\sum_{n\geq 1}^{\mathcal{R}}$, and then we have
@@
We give in section 2, a short introduction to Ramanujan’s summation of divergent series. Since we will see that, for $s\neq 1$, we have
@@
and $r_{m}$ are the rational numbers
$r_{m}=(-1)^{m+1}+\sum_{k=1}^{m}\frac{1+(-1)^{k-1}}{k}.$
We prove that the roots of $P_{m}$ are on the line $\{\Re(s)=\frac{1}{2}\}$. This is a consequence of the fact that the polynomials
$Q_{m}:t\mapsto\frac{1}{2}P_{m}(\frac{1}{2}+it)$
form a complete orthonormal system in $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. The expansion
@@
is an orthonormal expansion in the space $L^{2}\big(\mathbb{R},\frac{dt}{\cosh(\pi t)}\big)$. Also by a formal calculation, we get
$\frac{1}{2}\int_{-\infty}^{+\infty}\Big|\zeta(\frac{1}{2}+it)\Big|^{2}\frac{dt}{\cosh(\pi t)}=\log(2\pi)-\gamma-\frac{1}{2}.$
@@ 2.1 Ramanujan’s constant of a series
Let a function $f\in C^{1}(]0,+\infty[)$, then for every integer $n\geq 1$, we have the first order Euler-MacLaurin formula
$\sum_{k=1}^{n}f(k)=\int_{1}^{n}f(x)\,dx+\frac{1}{2}\big(f(n)+f(1)\big)+\int_{1}^{n}b(x)f^{\prime}(x)\ dx$}, (1)
where
b(x)=x[x]12 and [x] denote the integer part of x.b(x)=x-[x]-$\frac${1}{2}\ $\text${ and }\ [x]\ $\text${ denote the integer part of }\ x.
Let’s assume that the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\ dx$ is convergent. The previous formula gives
@@
We observe that $C(f)$ is also well defined for some divergent series. For example, we have immediately $C(1)=1/2,$ and applying (2) to the function $f:x\mapsto 1/x$, we have for any integer $n\geq 1$
@@
and we see that the Ramanujan’s constant of the series $\sum_{n\geq 1}\frac{1}{n}$ is the Euler constant :
$\gamma=\lim_{n\to+\infty}\Big(\sum_{k=1}^{n}\frac{1}{k}-\log n\Big).$
@@ 2.2 Plana’s formula
Let $f$ be an analytic function in the half-plane $\{\Re(z)>0\}$ such that there is a constant ${a}<2\pi$ and, for any interval $c,d]\subset]0,+\infty[$, a constant $M>0$ with
$\left|f(x+it)\right|\leq M\ e^{a|t|},\ \text{ for }x\in[c,d]\ \text{ and }t\in\mathbb{R}.$
@@
If we now compare Euler-MacLaurin formula (1) and Plana’s formula (3), then we get, for $n\geq 1$,
@@
Then if the integral $\int_{1}^{+\infty}b(x)f^{\prime}(x)\,dx$ is convergent and if
@@
Thus, in this case the Ramanujan’s constant $C(f)$, previously defined in (2), is also
@@ 2.3 The space $\mathcal{E}^{\pi}$
Let $\mathcal{E}$} the space of analytic functions $f$ in the half plane $\{\Re(z)>0\}$ such that there is some constants $\emph{a}<\pi$ and $\emph{b}>0$ with :
@@
Thus $f_{s}\in\mathcal{E}$ if and only if $\pi<\Im(s)<\pi$.
@@
If $f\in\mathcal{E}$ is such that $f(n)=0$ for all $n\;=\;1,\,2,\,\ldots$, then $f\equiv 0$.
Thus a function $f\in\mathcal{E}$ is only determined by the values $f(1),f(2),f(3),\dots$.
@@
The function $g\;:\;z\mapsto\sin(\pi z)$ is such that $g(n)=0$ for every integer $n\geq 1$}, but $g\notin\mathcal{E}$ since for $x\in\mathbb{R}$ and $t\in\mathbb{R}$
@@ 2.4 Ramanujan’s method of summation
Let $f$ and $g$ two functions verifying the hypothesis of theorem 1. By Carlson’s theorem we know that if $f\in\mathcal{E}$ and $g\in\mathcal{E}$ verify f(n)=g(n){f(n)=g(n)} for $n=1,2,3,\dots$, then for x>0{x>0} and $t\in\mathbb{R}$, we have
@@
For $f\in\mathcal{E}$, we define the Ramanujan’s summation of the series $\sum_{n\geq 1}f(n)$ by
$\sum_{n\geq 1}^{\mathcal{R}}f(n)=C(f)=\frac{f(1)}{2}+i\!\int_{0}^{+\infty}\frac{f(1+it)-f(1-it)}{e^{2\pi t}-1}\,\ dt\,.$ (6)
@@
Let $f\in\mathcal{E}$ then by Plana’s formula (3), we have for every integer $n\geq 1$
@@
and if in addition, the integral $\int_{1}^{+\infty}f(u)\ du$ is convergent then the series $\sum_{n\geq 1}f(n)$ is convergent. In this case (7) gives the relation
@@
Let $f:x\mapsto\frac{1}{x^{s}}$, then for $\Re(s)>1$, we get
@@ 2.5 Interpolation of partial sums
To evaluate Ramanujan’s summation of a divergent series $\sum_{n\geq 1}f(n)$, a simpler way than the formula (6) can be obtained using an interpolation function of the partial sums of this series. In chapter VI of his Notebooks, Ramanujan uses, without any definition, such a function $\varphi$ which satisfies
@@
a) Existence. Since $f\in\mathcal{E}$ then the function $\varphi_{f}$, defined for $\Re(z)>0$, by
@@
is also in $\mathcal{E}$. By (3) we have $\varphi_{f}(n)=\sum_{k=1}^{n}f(k)$ for every integer $n\geq 1$, thus the function $\varphi_{f}$ verifies
@@
Since the functions $z\mapsto\varphi_{f}(z+1)-\varphi_{f}(z)$ and $z\mapsto f(z+1)$ are in $\mathcal{E}$, then by Carlson’s theorem, we get
@@
The condition $\varphi_{f}(1)=f(1)$ is simply a consequence of (10) and definition (6).
b) Uniqueness. If $\varphi\in\mathcal{E}$ is such that
@@
thus $\varphi(n)=\sum_{k=1}^{n}f(k)=\varphi_{f}(n)$ for every integer $n\geq 1$. Since $\varphi$ and $\varphi_{f}$ are in the space $\mathcal{E}$, then by Carlson’s theorem we get $\varphi=\varphi_{f}.$
@@
The preceding proof asserts that if $f\in\mathcal{E}$ then the function $\varphi_{f}$ given by (10) satisfies
@@
Thus this function $\varphi_{f}$ is the unique function in $\mathcal{E}$ which interpolates the partial sums of the series $\sum_{n\geq 1}f(n)$. Note that, by (11), we have for $\Re(z)>0$
@@
Let the function $f:z\mapsto 1/z$. We will ses that $\varphi_{f}$ is related to digamma function which is defined, for $\Re(z)>0$, by
@@
The function $\Psi$ is analytic in $\{\Re(z)>0\}$, and by logarithmic differentiation of the classical product formula for $\Gamma$ (cf. [G-R]), we get for $\Re(z)>0$
$\Psi(z)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\big(\frac{1}{n}-\frac{1}{n+z}\big)=-\frac{1}{z}-\gamma+\sum_{n=1}^{+\infty}\frac{z}{n(n+z)}.$
Thus it is easily verified that $\Psi\in\mathcal{E}$ and that
$\Psi(z+2)\;-\;\Psi(z+1)=\frac{1}{z+1}.$
@@
Let $f\in\mathcal{E}$, and let $\varphi\in\mathcal{E}$ be a solution of
$\varphi(x+1)-\varphi(x)=f(x+1),\ \text{ for all }\ x>0,\ \text{ and }\ \varphi(1)=f(1).$
@@
$\int_{1}^{2}\varphi_{f}(x)\ dx\;\;=\int_{1}^{2}F(u)\ du\,+\,\frac{1}{2}\int_{1}^{2}f(x)\ dx\,+\sum_{n\geq 1}^{\mathcal{R}}f(n)-iD(f),$
@@
If we apply Plana’s formula (3) to the function $F$, in the case $n=2$, then we get
$-\big(\frac{1}{2}F(1)+\frac{1}{2}F(2)\big)+\!\int_{1}^{2}F(t)\ dt=i\int_{0}^{+\infty}\frac{F(2+it)-F(2-it)}{e^{2\pi t}-1}\ dt-i\int_{0}^{+\infty}\frac{F(1+it)-F(1-it)}{e^{2\pi t}-1}\ dt,$
that is
@@
Let $f\in\mathcal{E}$, then the function $x\mapsto f(x+1)$ belongs to $\mathcal{E}$ and we have the shift formula
$\sum_{n\geq 1}^{\mathcal{R}}f(n+1)\;\;=\sum_{n\geq 1}^{\mathcal{R}}f(n)-f(1)+\int_{0}^{1}f(x+1)\ dx.$ (12)
@@
Let $f:z\mapsto 1/z$. By example 3, we have $\varphi_{f}(x)=\Psi(x+1)+\gamma,$ thus we get
@@
Observe that, by the shift formula (12), we have
@@
Let $\log$ the principal determination of logarithm. Since, for any $x>0$, we have
$\log(\Gamma(x+2))-\log(\Gamma(x+1))=\log(x+1),$
@@
Let $f=\Psi$ and let $\varphi:x\mapsto x\Psi(x)-x+1$. We have $f(1)=\Psi(1)=\varphi(1)$ and
@@
Thus, by the previous theorem, we have
@@
then by Raabe’s formula, we get
$\sum_{n\geq 1}^{\mathcal{R}}\Psi(n)=-\log(\sqrt{2\pi})+\frac{1}{2}.$ (15)
@@
We now evaluate $\sum_{n\geq 1}^{\mathcal{R}}n^{k}$ for any $k$ integer $\geq 1$. Let the Bernoulli polynomials $B_{k}(x)$ be defined by $B_{0}(x)=1$ and, for $k\geq 1,$
$B^{\prime}_{k}(x)=kB_{k-1}(x)\ \text{ with }\ \int_{0}^{1}B_{k}(x)\ dx=0.$
The Bernoulli numbers are the numbers $B_{k}=B_{k}(0)$. For $k\geq 1$, we have $B_{k+1}(1)=B_{k+1}$. These polynomials verify (cf. [G-R])
@@
Let the function $f:x\mapsto x^{k}$ with $k$ integer $\geq 1$. Then by theorem 3, we obtain
@@
$\sum_{n\geq 1}^{\mathcal{R}}n^{0}=\frac{1}{2}=\zeta(0)+1.$
These formulae are extensions of formula (8) and are special cases the following general result.
@@
$\zeta(s)=\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}+\frac{1}{s-1}.$ (17)
@@
Let $s=\sigma+i\tau$ with $\sigma\in\mathbb{R},\ \tau\in\mathbb{R}$, and $z=x+it$ with $x>\varepsilon>0$, $t\in\mathbb{R}$. We have
@@
Thus for every $s\in\mathbb{C}$, the function $z\mapsto 1/{z^{s}}$ is in the space $\mathcal{E}$ and the sum $\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}$ is well defined. By (6), we have
$\sum_{n\geq 1}^{\mathcal{R}}\frac{1}{n^{s}}=\frac{1}{2}+i\int_{0}^{+\infty}\Big({\frac{1}{(1+it)^{s}}-\frac{1}{(1-it)^{s}}}\big)\frac{1}{e^{2\pi t}-1}\,\ dt,$